# In Search of a Hidden Curve

### v1

@@ 3.3. Variation 3: Spectral curves and opers
(3.24) $(E_{0},\phi)=\left(K_{C}^{\frac{1}{2}}\oplus K_{C}^{-{\frac{1}{2}}},\begin{bmatrix}&q\\ 1&\end{bmatrix}\right)$
@@
because $\phi$ is a matrix valued $1$-form. We note that $dw/dz=\xi_{wz}^{-2}=1/(cz+d)^{2}$. As calculated in [32, Section 3.1], the extension group ${\text{\rm{Ext}}}^{1}\big{(}K_{C}^{-{\frac{1}{2}}},K_{C}^{{\frac{1}{2}}}\big{)}$ is generated by the cohomology class $\frac{d}{dz}\xi_{wz}=c$, which is the image of the line bundle $K_{C}^{\frac{1}{2}}$ under the cohomology map
@@
The precise cancellation here makes us feel miraculous. ∎
Definitions of a few terminologies are due. Deligne’s $\hbar$-connection $\nabla^{\hbar}$ is a linear differential operator $\nabla^{\hbar}:E\longrightarrow E\otimes K_{C}[[\hbar]]$ on a vector bundle $E$ that satisfies
@@
$\nabla\big{|}_{{\mathcal{F}}^{i}}:{\mathcal{F}}^{i}\longrightarrow{\mathcal{F}}^{i-1}\otimes K_{C},\quad i=1,2,\dots,n,$
@@
¯$\overline{\nabla}:{\mathcal{F}}^{i}\big{/}{\mathcal{F}}^{i+1}\overset{\sim}{\longrightarrow}\big{(}{\mathcal{F}}^{i-1}\big{/}{\mathcal{F}}^{i}\big{)}\otimes K_{C},\quad i=1,2,\dots,n-1,$
@@
The discovery of [32, Section 3.1] is a full generalization of Proposition 3.4 to the case of the Hitchin section on ${\mathcal{M}}\big{(}C,SL_{n}({\mathbb{C}})\big{)}$.
@@
4. Finale Incompiuto: The Geometry that knows the irrationality of $\zeta(3)$
@@ 4.1. The Mystery Formula
(4.1) $(-1)^{n}(n!)^{2}$ $\sum_{\begin{subarray}{c}\ell+m=n\\ \ell,m\geq 0\end{subarray}}\frac{(2\ell+m)!(\ell+2m)!}{(\ell!)^{5}(m!)^{5}}\bigg{(}1+(m-\ell)\big{(}H_{2\ell+m}+2H_{m+2\ell}-5H_{m}\big{)}\bigg{)}$
@@
Apéry’s astonishing discovery of [3] announced in 1978, that proves the irrationality of a special value of the Riemann zeta function, $\zeta(3)$, caused a huge sensation (see [92]). It was followed by an almost immediate, and unexpected, geometric surprise [8, 9] that the generating function of the key integer sequence $A_{n}$ of Apéry, the second line of (4.1), solves a Picard-Fuchs differential equation associated with a particular family of K3 surfaces. Since last decade, there has been a renewed sensation [43, 44, 45, 64, 96] on this topic. This time it is due to the identification that the Borel-Laplace transform of the same generating function is identical to the generating function of the genus 00, $1$-marked point, degree $d\geq 2$ Gromov-Witten invariants of a Fano $3$-fold described below, after adjusting the exponential factor $e^{5t}$. This is the content of Mystery Formula (4.1).
@@
We consider a rank $3$ vector bundle ${\mathcal{E}}=({\mathcal{U}}^{*}\otimes\det{\mathcal{U}}^{*})\oplus\det{\mathcal{U}}^{*}$ defined on the $6$-simensional Grassmannian $Gr(2,5)$, where ${\mathcal{U}}$ is the universal bundle. Take a generic section $\sigma\in H^{0}\big{(}Gr(2,5),{\mathcal{E}})$ of ${\mathcal{E}}$. The Fano $3$-fold $V_{12}$ is the zero locus $[\sigma]_{0}$ of this section. The genericness condition assures smoothness of the zero locus.
@@
Let $X$ denote a non-singular model for the Fano $3$-fold $V_{12}$. The degree $d\geq 2$, type $(g,n)=(0,1)$ Gromov-Witten invariant of $X$ is defined by
(4.2) $\gamma_{d}:=\int_{{\overline{\mathcal{M}}}_{0,1}(X)}ev^{*}(pt)\psi^{d-2},$
@@
  • ¯${\overline{\mathcal{M}}}_{0,1}(X)$ is the moduli stack of holomorphic maps $f:({\mathbb{P}}^{1},\infty)\longrightarrow X$,
  • $ev:{\overline{\mathcal{M}}}_{0,1}(X)\owns f\longmapsto f(\infty)\in X$ is the evaluation map,
  • $ev^{*}:H^{\bullet}(X,{\mathbb{Q}})\longrightarrow H^{\bullet}\big{(}{\overline{\mathcal{M}}}_{0,1}(X),{\mathbb{Q}})$ is the induced cohomology map, and
  • $\psi=c_{1}({\mathbb{L}})$ is the first Chern class of the tautological line bundle $\pi:{\mathbb{L}}\longrightarrow{\overline{\mathcal{M}}}_{0,1}(X)$ whose fiber at an object $[f]\in{\overline{\mathcal{M}}}_{0,1}(X)$ is $T^{*}_{f(\infty)}f({\mathbb{P}}^{1})$.
To make the formula simpler, we choose $\gamma_{0}=1,\gamma_{1}=0$. Then the generating function of these type $(0,1)$ Gromov-Witten invariants is given by
(4.3) $GW_{0,1}(X):=\sum_{d=0}^{\infty}\gamma_{d}\;t^{d}=\;e^{-5t}$ $\sum_{n=0}^{\infty}(-1)^{n}n!\sum_{\begin{subarray}{c}\ell+m=n\\ \ell,m\geq 0\end{subarray}}\frac{(2\ell+m)!(\ell+2m)!}{(\ell!)^{5}(m!)^{5}}$
$\times$ $\bigg{(}1+(m-\ell)\big{(}H_{2\ell+m}+2H_{m+2\ell}-5H_{m}\big{)}\bigg{)}\;t^{n}.$
We recall that topological recursion for simple Hurwitz numbers and higher genus Catalan numbers, as we have seen in earlier sections, are uniformly formulated on the corresponding spectral curves, and the spectral curves are exactly the generating functions of the $(g,n)=(0,1)$-invariants in each case. The role of topological recursion is to produce all $(g,n)$ invariants from the spectral curve through a universal recursive procedure.
@@
Nothing precise is known at this moment. The Quantum Differential Equation (see [52]) for the Fano variety $V_{12}$ is identified in [18] (see below, Theorem 4.17). The quantum curve in question is an $\hbar$-deformation family of QDE. This point relates to the author’s work [27] on Gaiotto conjecture [41] discussed in Section 3.
@@ 4.2. A quick review of Apéry’s ideas
Partial sums of $\zeta(3)=\sum_{m=1}^{\infty}\frac{1}{m^{3}}$ are appearing here.
@@
In particular, both $A_{n}$ and $B_{n}$ grow exponentially fast, $\sim n^{-\frac{3}{2}}C^{n}$, as $n\rightarrow\infty$.
@@
$10<\left(\frac{xyz(1-x)(1-y)(1-z)}{1-z+xyz}\right)^{n}<$ 1
@@
We know $A_{n}$ and $B_{n}$ grow exponentially fast, while the linear combination $A_{n}\zeta(3)-B_{n}$ converges to 00. It is also a solution of the recursion (4.4). Therefore, it decreases exponentially fast!
(4.12) $A_{n}\zeta(3)-B_{n}\sim{\text{\rm{const}}}\;n^{-\frac{3}{2}}C^{-n}\Longrightarrow\left|\zeta(3)-\frac{B_{n}}{A_{n}}\right|\sim{\text{\rm{const}}}\;C^{-2n},\quad n\rightarrow\infty.$
@@
$\left|r-\frac{p}{q}\right|<\frac{1}{q^{1+\epsilon}}.$
@@
$\frac{1}{q^{1+\epsilon}}>\left|r-\frac{p}{q}\right|=\left|\frac{a}{b}-\frac{p}{q}\right|=\left|\frac{aq-bp}{bq}\right|\geq\frac{1}{bq},\quad\text{hence}\quad 0<q<b^{1/\epsilon}.$
@@
He found that $d_{n}^{3}B_{n}\in{\mathbb{Z}}$, which follows from (4.6). The recursion (4.4) suggests that the denominator of a solution $u_{n}$ starting with integral initial values $u_{0},u_{1}\in{\mathbb{Z}}$ is about $(n!)^{3}$. The denominator of $B_{30}$ is $\sim 1.9\times 10^{33}$, while LCM [${\text{\it{LCM\,}}}[1,\dots,30]^{3}$ divided by this denominator is $6630$. So the above estimate is quite accurate.
@@
$\left|\zeta(3)-\frac{B_{n}}{A_{n}}\right|\sim{\text{\rm{const}}}\;C^{-2n}<\frac{1}{\left(n^{-\frac{3}{2}}d_{n}^{3}C^{n}\right)^{1+\epsilon}},\qquad n\gg 1.$
@@ 4.3. The differential equations behind the scene
(4.15) $\begin{cases}PA(t)=0\\ Q\alpha(t)=0,\end{cases}\qquad\begin{cases}PB(t)=6t\\ Q\beta(t)=6t.\end{cases}$
@@ Appendix A The Lagrange Inversion Formula
[…]
@@
[…]

### v2

@@ 3.3. Spectral curves and opers
[…]
@@
because $\phi$ is a matrix valued $1$-form. We note that $dw/dz=\xi_{wz}^{-2}=1/(cz+d)^{2}$. As calculated in [33, Section 3.1], the extension group ${\text{\rm{Ext}}}^{1}\big(K_{C}^{-{\frac{1}{2}}},K_{C}^{{\frac{1}{2}}}\big)$ is generated by the cohomology class $\frac{d}{dz}\xi_{wz}=c$, which is the image of the line bundle $K_{C}^{\frac{1}{2}}$ under the cohomology map
@@
The precise cancellation here makes us feel miraculous. Definitions of a few terminologies are due here. Deligne’s $\hbar$-connection $\nabla^{\hbar}$ is a linear differential operator $\nabla^{\hbar}:E\longrightarrow E\otimes K_{C}[[\hbar]]$ on a vector bundle $E$ that satisfies
@@
$\nabla\big|_{{\mathcal{F}}^{i}}:{\mathcal{F}}^{i}\longrightarrow{\mathcal{F}}^{i-1}\otimes K_{C},\qquad i=1,2,\dots,n,$
@@
$\overline{\nabla}:{\mathcal{F}}^{i}\big/{\mathcal{F}}^{i+1}\overset{\sim}{\longrightarrow}\big({\mathcal{F}}^{i-1}\big/{\mathcal{F}}^{i}\big)\otimes K_{C},\qquad i=1,2,\dots,n-1,$
@@
The discovery of [33, Section 3.1] is a full generalization of Proposition 3.4 to the case of the Hitchin section on ${\mathcal{M}}\big(C,SL_{n}({\mathbb{C}})\big)$.
@@
4. In search of the geometry that knows the irrationality of $\zeta(3)$
@@ 4.1. The Mystery Formula
(4.1) $(-1)^{n}(n!)^{2}$ $\sum_{\begin{subarray}{c}\ell+m=n\\ \ell,m\geq 0\end{subarray}}\frac{(2\ell+m)!(\ell+2m)!}{(\ell!)^{5}(m!)^{5}}\bigg(1+(m-\ell)\big(H_{2\ell+m}+2H_{\ell+2m}-5H_{m}\big)\bigg)$
@@
Apéry’s astonishing discovery of [4] announced in 1978, that proves the irrationality of a special value $\zeta(3)$ of the Riemann zeta function, caused a huge sensation (see [95]). It was followed by an almost immediate, and unexpected, geometric surprise [9, 10] that the generating function of the key integer sequence $A_{n}$ of Apéry, the second line of (4.1), solves a Picard-Fuchs differential equation associated with a particular family of K3 surfaces. Since last decade, there has been a renewed sensation [44, 46, 47, 68, 99] on this topic. This time it is due to the identification that the Borel-Laplace transform of the same generating function is identical to the generating function of the genus 0, 11-marked point, degree $d\geq 2$ Gromov-Witten invariants of a Fano $3$-fold described below, after adjusting the exponential factor $e^{5t}$. This is the content of the Mystery Formula (4.1).
@@
We consider a rank $3$ vector bundle ${\mathcal{E}}=({\mathcal{U}}^{*}\otimes\det{\mathcal{U}}^{*})\oplus\det{\mathcal{U}}^{*}$ defined on the $6$-simensional Grassmannian $Gr(2,5)$, where $\mathcal{U}$} is the universal bundle. Take a generic section $\sigma\in H^{0}\big(Gr(2,5),{\mathcal{E}})$ of $\mathcal{E}$}. The Fano $3$-fold $V_{12}$ is the zero locus $\sigma]_{0}$ of this section. The genericness condition assures smoothness of the zero locus.
@@
Let $X$ be a non-singular model for the Fano $3$-fold $V_{12}$, and $K_{X}$ its anticanonical divisor. Define the anticanonical degree $d\geq 2$, type $(g,n)=(0,1)$, Gromov-Witten invariant of $X$ by
(4.2) $\gamma_{d}:=\sum_{\begin{subarray}{c}\beta\in H_{2}(X,{\mathbb{Z}})\\ d={\langle}\beta,-K_{X}{\rangle}\end{subarray}}\int_{{\overline{\mathcal{M}}}_{0,1}(X,\beta)}ev^{*}(pt)\psi^{d-2},$
@@
  • ${\overline{\mathcal{M}}}_{0,1}(X,\beta)$ is the moduli stack of holomorphic maps $f:(C,\infty)\longrightarrow X$ from a genus 0 curve $C$ with a marked point $\infty\in C$ to $X$ such that the homology class of $f(C)$ is equal to $\beta$ in $H_{2}(X,{\mathbb{Z}})$;
  • $ev:{\overline{\mathcal{M}}}_{0,1}(X,\beta)\owns f\longmapsto f(\infty)\in X$ is the evaluation map;
  • $ev^{*}:H^{\bullet}(X,{\mathbb{Q}})\longrightarrow H^{\bullet}\big({\overline{\mathcal{M}}}_{0,1}(X,\beta),{\mathbb{Q}})$ is the induced cohomology map; and
  • $\psi=c_{1}({\mathbb{L}})$ is the first Chern class of the tautological line bundle $\pi:{\mathbb{L}}\longrightarrow{\overline{\mathcal{M}}}_{0,1}(X,\beta)$ whose fiber at $f\in{\overline{\mathcal{M}}}_{0,1}(X,\beta)$ is $T^{*}_{f(\infty)}f(C)$.
Then the generating function of these type $(0,1)$ Gromov-Witten invariants is given by
(4.3) $GW_{0,1}(X)=\sum_{d=0}^{\infty}\gamma_{d}\;t^{d}=\;e^{-5t}$ $\sum_{n=0}^{\infty}(-1)^{n}n!\sum_{\begin{subarray}{c}\ell+m=n\\ \ell,m\geq 0\end{subarray}}\frac{(2\ell+m)!(\ell+2m)!}{(\ell!)^{5}(m!)^{5}}$
$\times$ $\bigg(1+(m-\ell)\big(H_{2\ell+m}+2H_{\ell+2m}-5H_{m}\big)\bigg)\;t^{n}.$
Remark 4.3.
From the expected dimension $\dim{\overline{\mathcal{M}}}_{0,1}(X,\beta)=d+1$, we normalize $\gamma_{0}=1$ and $\gamma_{1}=0$. Only $\gamma_{2}$ has a direct geometric counting interpretation. The exponential factor is chosen to assure $\gamma_{1}=0$.
We recall that topological recursion for simple Hurwitz numbers and higher genus Catalan numbers, as we have seen in earlier sections, are uniformly formulated on the corresponding spectral curves, and the spectral curves are exactly the generating functions of the $(g,n)=(0,1)$-invariants. The role of topological recursion is to produce all $(g,n)$ invariants from the spectral curve through a universal recursive procedure.
@@
Nothing precise is known at this moment. The Quantum Differential Equation (see [55]) for the Fano variety $V_{12}$ is identified in [19] (see below, Theorem 4.17). The quantum curve in question is an $\hbar$-deformation family of QDE. This point relates to the work [28] on Gaiotto conjecture [42] discussed in Section 3.
@@ 4.2. A quick review of Apéry’s ideas
Partial sums of $\zeta(3)=\sum_{m=1}^{\infty}\frac{1}{m^{3}}$ are appearing here. Notice that $\{A_{n}\}$ is a natural basis for the space of solutions because it is unique, up to a constant multiplication, that satisfies (4.4) for all $n\geq 0$.
@@
In particular, both $A_{n}$ and $B_{n}$ grow exponentially fast, $\sim n^{-\frac{3}{2}}C^{n}$, as $n\rightarrow\infty$. Since the space of solutions has dimension $2$, there is a unique solution, again up to a constant multiple, that decays exponentially. This means there exists a unique constant $\zeta\in{\mathbb{R}}$ such that
$\lim_{n\rightarrow\infty}\big(A_{n}\zeta-B_{n}\big)=0.$
Since $\{A_{n}\}$ is unique up to constant multiple, and $\{B_{n}\}$ is canonical up to constant multiple by choosing $B_{0}=0$ to make it linearly independent with $\{A_{n}\}$, the sequence (4.4) knows this constant $\zeta$.
@@
$0<\left(\frac{xyz(1-x)(1-y)(1-z)}{1-z+xyz}\right)^{n}<1$
@@
The mysterious constant $\zeta$ is indeed $\zeta(3)$. We know $A_{n}$ and $B_{n}$ grow exponentially fast, while the linear combination $A_{n}\zeta(3)-B_{n}$ converges to 0. It is also a solution of the recursion (4.4). Therefore, it decreases exponentially fast!
(4.12) $A_{n}\zeta(3)-B_{n}\sim{\text{\rm{const}}}\;n^{-\frac{3}{2}}C^{-n}\Longrightarrow\left|\zeta(3)-\frac{B_{n}}{A_{n}}\right|\sim{\text{\rm{const}}}\;C^{-2n},\qquad n\rightarrow\infty.$
@@
$0<\left|r-\frac{p}{q}\right|<\frac{1}{q^{1+\epsilon}}.$
@@
$\frac{1}{q^{1+\epsilon}}>\left|r-\frac{p}{q}\right|=\left|\frac{a}{b}-\frac{p}{q}\right|=\left|\frac{aq-bp}{bq}\right|\geq\frac{1}{bq},\qquad\text{hence}\qquad 0<q<b^{1/\epsilon}.$
@@
He found that $d_{n}^{3}B_{n}\in{\mathbb{Z}}$, which follows from (4.6). The recursion (4.4) suggests that the denominator of a solution $u_{n}$ starting with integral initial values $u_{0},u_{1}\in{\mathbb{Z}}$ is about $(n!)^{3}$. The denominator of $B_{30}$ is $\sim 1.9\times 10^{33}$, while ${\text{\it{LCM\,}}}[1,\dots,30]^{3}$ divided by this denominator is $6630$. So the above estimate is quite accurate.
@@
$\left|\zeta(3)-\frac{B_{n}}{A_{n}}\right|\sim{\text{\rm{const}}}\;C^{-2n}<\frac{1}{\left(n^{-\frac{3}{2}}d_{n}^{3}C^{n}\right)^{1+\epsilon}},\hskip 23.49976ptn\gg 1.$
@@ 4.3. The differential equations behind the scene
(4.15) $\begin{cases}PA(t)=0\\ Q\alpha(t)=0,\end{cases}\hskip 23.49976pt\begin{cases}PB(t)=6t\\ Q\beta(t)=6t.\end{cases}$
@@
5. Concluding Remark
Our observation is that the generating function of all $g=0$, $n=1$ invariants for a particular counting problem, such as Catalan numbers or tree counts as discussed earlier, determines a spectral curve. In terms of complex analysis over the spectral curve, we can generate all $(g,n)$ invariants through a recursive mechanism. The $(0,1)$ invariants themselves satisfy a recursion, which is used to determine the analytic formula for the spectral curve. And in terms of the rightly chosen coordinate of the spectral curve, the generating functions of the $(g,n)$ invariants turn out to be polynomials. The asymptotic behavior of these polynomials then recover the Witten-Kontsevich formula, Virasoro constraint conditions, and particular counting problems including the $\lambda_{g}$-formula and the Euler characteristic of ${\mathcal{M}}_{g,n}$.
The fact that such mechanism exists makes the starting counting problem very special. The method presented in these lectures are not meant to be a subject for generalization. Rather, it tells us how special these problems are, and this explains why they have deep connections with integrable systems, representation of Virasoro algebras, moduli spaces of algebraic curves, and so on.
Apéry’s recursions are in the sense extremely special. Again, they are not a subject for generalization. We are led to wondering how and why they are so special. Once we understand these, we may be able to transplant the geometric situation to other equally special cases, and may obtain a new arithmetic result.
The Apéry recursions for $\zeta(2)$ and $\zeta(3)$ are indeed recursions for the $(0,1)$ case of Gromov-Witten invariants of very special target manifolds. As such, the geometric origin of these recursions come from degenerations of rational curves on these particular target spaces. In this way we understand the origin of the Apéry recursions from symplectic geometry. Its mirror partner is the Gauss-Manin connection found in this context, which generate infinite sequence of periods. The speed of the asymptotic behavior of these sequences translates into the irrationality of $\zeta(2)$ and $\zeta(3)$.
It reminds us of the new proof of the Lawrence conjecture [62] due to Habiro, on the integrality of certain $3$-manifold invariants over the ring generated by a root of unity using his powerful universal invariants [50]. When considered for more general link invariants, one understands the amazing symplectic nature of the counting problem (see for example, [3]). The role that the universal invariants play, in particular, in the context of arithmeticity [45], suggests the parallel mirror symmetric point of view in these developments.
${\boxed{\text{Periods}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\boxed{\text{Symplectic Counting}}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\boxed{\text{Mirror Symmetry}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\boxed{\text{Connections on Families}}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\boxed{\text{Quantization}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@ Appendix A The Lagrange Inversion Formula
$y-b$ $=\int_{b}^{y}1\cdot ds=\int_{b}^{y}\left(\frac{1}{2\pi i}\oint\frac{f^{\prime}(s)dt}{f(t)-f(s)}\right)ds$
$=\int_{b}^{y}\left(\frac{1}{2\pi i}\oint\frac{f^{\prime}(s)dt}{\big(f(t)-a\big)-\big(f(s)-a\big)}\right)ds$
$=\frac{1}{2\pi i}\int_{b}^{y}\oint\frac{\frac{f^{\prime}(s)}{f(t)-a}}{1-\frac{f(s)-a}{f(t)-a}}dtds$
$=\frac{1}{2\pi i}\int_{b}^{y}\sum_{n=0}^{\infty}\oint\frac{f^{\prime}(s)}{f(t)-a}\left(\frac{f(s)-a}{f(t)-a}\right)^{n}dsdt$
$=\frac{1}{2\pi i}\int_{f(b)}^{f(y)}\sum_{n=0}^{\infty}\oint\frac{1}{f(t)-a}\left(\frac{f(s)-a}{f(t)-a}\right)^{n}df(s)dt$
$=\frac{1}{2\pi i}\sum_{n=0}^{\infty}\oint\frac{1}{\big(f(t)-a\big)^{n+1}}\cdot\frac{\big(f(y)-a\big)^{n+1}}{n+1}dt$
@@
$=\frac{1}{2\pi i}\sum_{k=1}^{\infty}\oint\frac{1}{(y-b)^{k}}\left(\frac{y-b}{f(y)-a}\right)^{k}dy\cdot\frac{(x-a)^{k}}{k}$